OpenStudy (anonymous):

Huilan's age is two times Thomas's age. The sum of their ages is 21 . What is Thomas's age?

OpenStudy (beccaboo333):

Alright, do you have any idea how to do this?

OpenStudy (beccaboo333):

Sum=added up total. So we know that, if Huilan's age is twice what Thomas's age is.. What could that possibly be? Any ideas?

OpenStudy (anonymous):

i'm not sure, I'm not too good with these type of problems, but I just need the answer to this one to finish my work

OpenStudy (beccaboo333):

Well, we can't just give straight up answers on here.

OpenStudy (beccaboo333):

I'll work through it with you.

OpenStudy (anonymous):

I know I just simply don't understand this, I'm asking my teacher tomorrow but I just dont get it :/

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I wanna say Hulian is about 4 or 5

OpenStudy (anonymous):

*Thomas i meant

OpenStudy (beccaboo333):

No. That would add up to either 12 or 15 . Not 21. >.>

OpenStudy (anonymous):

I know, I'm completely lost

OpenStudy (anonymous):

7

OpenStudy (anonymous):

Oh I get it

OpenStudy (anonymous):

Thomas is 7 and Hulian is 14? 7 + 14 = 21?

OpenStudy (anonymous):

Thanks, can you help me with one or 2 more problems like this?

OpenStudy (beccaboo333):

Close this and ask it as a new :3

OpenStudy (anonymous):

ok

OpenStudy (beccaboo333):

@iambatman will correct my mistakes or explain it better for me xD

OpenStudy (anonymous):

H = T + 2 H + T = 21 (T+2) + T = 21 2T = 19 T = 9.5

OpenStudy (anonymous):

Thomas is 9 and half years old :d